# DWTNeRF: Boosting Few-shot Neural Radiance Fields via Discrete Wavelet Transform

### v2

@@ 5.2 Qualitative Results
Method Venue PSNR ()($\uparrow$)( ↑ ) SSIM ()($\uparrow$)( ↑ ) LPIPS ()($\downarrow$)( ↓ )
@@ 5.3 Quantitative Results
On LLFF We compared DWTNeRF against recent few-shot works published at top venues. Regularization methods include InfoNeRF [16], FreeNeRF [39] and CombiNeRF [2]. Prior methods include DietNeRF [11], RegNeRF [23], DiffusioNeRF [37] and GeCoNeRF [17]. Model method includes mi-MLP [46]. Table 1 shows the quantitative results. Methods with the symbol $\dagger$ were reproduced on our machine, keeping all default settings. Unavailable metrics are denoted “--”. The best-performing metrics under each few-shot condition are bolded. In the 2-shot case, DWTNeRF outperforms CombiNeRF by 1.6% in PSNR, 0.84% in SSIM and 4.17% in LPIPS. The effects of our DWTNeRF are still strongly seen with 3-shot, but generally fade with higher views, suggesting its high compatibility with extreme few-shot conditions.
On NeRF-Synthetic Table 2 shows the quantitative results. In the 4-shot case, DWTNeRF outperforms CombiNeRF by 0.5% in PSNR, 0.1% in SSIM and 0.9% in LPIPS.
@@ 5.4 Ablation Studies
On main modules Those are the DW loss and the cross-banch interactions. Table 3 shows the effects of each module. The second row removes the frequency masking from CombiNeRF, given its incompatibility with INGP. This provides improvements in all metrics. The third row is our DWTNeRF, but with the DW loss only. Starting from the fourth row, we examine the effects of using multi-head attention in cross-branch interactions. The fourth row is DWTNeRF with only input-level attention. The fifth row only has output-level attention. The sixth row combines both levels. This produces stronger results over the previous two. The final row combines both DW loss and cross-branch interactions, which provides the strongest results.
@@
Table 4: Comparison against prior and model methods, LLFF dataset [20]
@@
CombiNeRF (baseline) $\dagger$ R 20.12 0.676 0.197
DietNeRF (+full cb.) $\dagger$ M, P 20.29 0.671 0.194
SPARF (+full cb.) $\dagger$ M, P 19.28 0.618 0.261
DWTNeRF (+full cb. + dw.) M, R 20.38 0.677 0.193
@@
On prior methods We implemented two prior-based approaches on top of CombiNeRF: DietNeRF’s semantic consistency (SC) loss [11] and SPARF’s multi-view (MV) correspondence loss [33]. We simply replaced the DW loss at Algorithm 1 with these losses. Table 4 (upper half) shows the results. “R”, “P”, “M” stand for regularization-, prior- and model-based, respectively. Our DW loss outperforms SC loss, but the DWT is much more lightweight than computing feature embeddings from a CLIP model. This suggests frequency supervision can be as important as semantic supervision. Performance worsens with MV loss, suggesting the multi-view correspondences are not a useful prior for INGP. For more implementation details, please refer to Sections 9 and 10 of Appendix.
On model methods We implemented two model-based approaches on top of CombiNeRF: “Residual Connections” (Figure 6a) and “Element-wise Cross-branch Interactions” (Figure 6b) - both introduced by mi-MLP [46]. Table 4 (lower half) shows the results. The second row introduces residuals. The third row is the same, but we added one MLP layer, hoping to better highlight the residuals’ effects. However, this degraded performances noticeably, highlighting the vulnerability of INGP to architectural changes. The fourth row shows mi-MLP’s cross-branch interactions. Only our version of interactions is competitive with the baseline. For more implementation details, please refer to Section 13 of Appendix.
Figure 6: mi-MLP’s [46] model methods: “Residual Connections” (a) and “Element-wise Cross-branch Interactions” (b)
On the choice of wavelets We compared the effects of different Daubechies wavelets [6] on the DW loss. We started with the Haar ($n=1$) and then higher-order wavelets, up to the db3 ($n=3$). Generally, higher-order wavelets have more supports, and spreads the signal energy more evenly across sub-bands. Section 8 of Appendix shows the exact wavelet coefficients. Figure 7 compares scene decomposition using the Haar versus db2 ($n=2$). Compared to the ground-truths, the rendered sub-bands are much noisier, highlighting the purpose of the DW loss. The db2 provides finer details. However, table 5 shows no significant differences between the wavelets. This confirms our intuition that low-frequency supervision is sufficient.
@@
Figure 7: Decomposition using the Haar (a & b) and db2 (c). We only show the LL (top) and LH (bottom) sub-bands. Column a) is the ground-truth LL, much less noiser than rendered LL at column b). Column c) uses higher-order wavelet, and shows finer details.
@@ 6 Conclusion
Limitations & Future Works Firstly, the DWT does not disentangle high-frequency components. For example, besides fine structural details, high frequencies can come from specularity and other light-dependent effects. Future works might explore scene decomposition at such further levels. Secondly, since the DW loss is used as an auxiliary loss alongside the photometric MSE loss, which compares pixel-wise differences regardless of frequencies, we believe some high-frequency overfitting still exists. Our DW loss prioritizes low frequencies, but not eliminate high frequencies in early training. Future works could tackle this problem, but interfering with the photometric loss can be tricky. We can also investigate the applicability of our methods to other representations, like SDF [3, 18] or 3DGS [15].
@@ 7 More Visualizations
We provide more qualitative results on the 3-shot LLFF and 4-shot NeRF-Synthetic benchmarks at Figures 13 and 12. There are 6 scenes in total. In each scene, we would like to direct attention towards the regions enclosed by the red box, which highlights the effects of cross-branch interactions (b $\rightarrow$ c) and DW loss (c $\rightarrow$ d). They provide finer details, preserve better structures and reduce random “hallucinated” points in 3D space.
@@ 8 Wavelet Coefficients
Table 6 shows the coefficients $\ell_{k}$ of the 1D low-pass filters $\bm{\ell}$ that are used in our experiments. The corresponding 2D low-pass matrices $\mathbf{L}$ are constructed with shifting rows of $\bm{\ell}$:
@@
and similarly for how $\mathbf{H}$ is constructed with the 1D high-pass filter $\bm{h}$. We experimented with orthogonal Daubechies wavelets [30], where $\bm{h}$ is derived trivially from $\bm{\ell}$ according to the “Alternating Flip” condition:
@@
where $k$ denotes a single index of the 1D filter with number of coefficients $N$. As a numerical example, suppose that $\bm{\ell}=[\ell_{0},\ell_{1},\ell_{2},\ell_{3}]$, the corresponding $\mathbf{h}$ that meets “Alternating Flip” condition is $\bm{h}=[\ell_{3},-\ell_{2},\ell_{1},-\ell_{0}]$. In summary, while the DW loss can technically be classified as a prior method, we classify it as a regularization method as it needs no more priors than the 1D low-pass filter $\bm{\ell}$. Besides its simplicity, the DW loss outperforms methods based on pre-trained weights like DietNeRF’s [11] semantic consistency loss and SPARF’s [33] multi-view correspondence loss - two highly popular representatives of prior methods. We discuss those methods in the next two sections.
@@
Figure 8: Frequency-based scene decompositions with different wavelets. From top to bottom: the LL, LH, HL and HH sub-bands. The first column (a) shows the sub-bands of ground-truth views using the Haar. The second column (b) is the same, but for rendered views, which are much noisier. The rest show sub-bands of rendered views using the db2 (c) and db3 (d). Higher-order wavelets provide more fine-grained details, but our experiments showed low-frequency supervision with coarse details is sufficient.
9 Frequency vs. semantic supervision
The semantic consistency (SC) loss [11] minimizes discrepancies in embeddings of rendered and ground-truth views:
$\begin{split}\mathcal{L}_{SC}=\frac{\lambda_{SC}}{2}||\phi[\mathbf{I}_{gt}(\mathbf{p})]-\phi[\mathbf{I}(\mathbf{\hat{p}})]||^{2}_{2}\end{split}$
where ϕ[]italic-$\phi[\cdot]$ represents the image embeddings. $\mathbf{I}(\mathbf{\hat{p}})$ and $\mathbf{I}_{gt}(\mathbf{p})$ are the rendered and ground-truth views sampled from an interpolated novel view $\mathbf{\hat{p}}$ and nearest known view $\mathbf{p}$, respectively. The intuition is that, an object is what it is, regardless of the viewing direction. This is the core principle of the semantic supervision, which encourages high-level semantic information to be consistent across views.
In practice, DietNeRF uses the CLIP [CLIP] embeddings, which are normalized. Since ϕ[]italic-$\phi[\cdot]$ is a unit vector, the semantic loss reduces to a cosine similarity weighted by the balancing term $\lambda_{SC}$:
$\begin{split}\mathcal{L}_{SC}=\lambda_{SC}\phi[\mathbf{I}_{gt}(\mathbf{p})]^{T}\phi[\mathbf{I}(\mathbf{\hat{p}})]\end{split}$
Implementation Details To compare the DW and SC losses fairly, we simply replaced the DW loss with the SC loss in Algorithm 1, and kept all other hyperparameters intact, except that the weight $\lambda_{SC}$ was set to 0.01 and the interval $K_{SC}$ was set to 30. Note that $K_{SC}>K_{DW}=10$ to make the training feasible, as calculating CLIP embeddings is computationally intensive. Similar to DietNeRF, to calculate the SC loss, we downsampled the known views to only 20% of the original resolution, and sampled rays across the downsampled image grid (so unlike DW loss, there was no use of patch-based rendering). The downsampling allows the rendering of novel views $\mathbf{I}(\mathbf{\hat{p}})$ to be computationally feasible. Furthermore, to sample a novel pose $\mathbf{\hat{p}}$, we sampled a pair of known views $\{\mathbf{p}_{i},\mathbf{p}_{j}\}$, and a weighting factor $\alpha\sim\mathcal{U}(0,1)$. Then, $\mathbf{\hat{p}}=\alpha\mathbf{p}_{i}+(1-\alpha)\mathbf{p}_{j}$. If $\alpha>0.5$, then $\mathbf{p}_{i}$ is the ground-truth pose and vice versa.
As presented in Table 4, the DW loss outperforms the SC loss with a comfortable margin, but requires no use of pre-trained weights. Given that semantic supervision is a popular method (e.g. in DietNeRF [11], GeCoNeRF [17], CaesarNeRF [CaesarNeRF]), we hope our results encourage fellow researchers to explore frequency supervision, a new kind of supervision that might be equally effective but absolutely less computationally demanding.
10 Frequency vs. multi-view supervision
Figure 9: Pixel matches predicted by a trained OmniGlue [OmniGlue], filtered by confidences (“fern” scene). There are a few incorrect matches, possibly reducing the effects of correspondence methods.
We additionally considered the multi-view supervision, utilized by the works of SPARF [33], CorresNeRF [CorresNeRF], ConsistentNeRF [ConsistentNeRF], GeCoNeRF [17], SGCNeRF [SGCNeRF], P2NeRF [P2NeRF], etc. For comparison, we implemented SPARF’s multi-view (MV) correspondence loss, which minimizes errors of pixel matches’ projections between known views:
$\mathcal{L}_{MV}=\lambda_{MV}\textbf{w}_{\mathbf{p}_{i},\mathbf{p}_{j}}f_{huber}[\mathbf{x}({\mathbf{p}_{j}})-\mathbf{x}_{gt}({\mathbf{p}_{j}})]$
where $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ represent two known poses. A trained feature matching model predicts the pixel matches between the two, $\{\mathbf{x}_{gt}({\mathbf{p}_{i}}),\mathbf{x}_{gt}({\mathbf{p}_{j}})\}$. We project the known matches in pose $\mathbf{p}_{i}$ to the 3D space, and then back onto another view at the pose $\mathbf{p}_{j}$. This gives the predicted matches $\mathbf{x}({\mathbf{p}_{j}})$, which should be the same as $\mathbf{x}_{gt}({\mathbf{p}_{j}})$ for a perfectly geometrically correct solution [multiview-book]. This is the multi-view correspondence principle. This loss constrains the implicit 3D model to be geometrically accurate. $f_{huber}$ is the Huber loss function [huber-book], weighted by the matches confidences $\textbf{w}_{\mathbf{p}_{i},\mathbf{p}_{j}}$ and the balancing term $\lambda_{MV}$.
Attempting to re-produce the best version of the MV loss, we supplemented it with ideas from follow-up works of SPARF. Firstly, we implemented CorresNeRF’s matches augmentation strategy [CorresNeRF], which swaps the target and query images to produce more matches. This is in line with SPARF’s observation that the training benefits from an increasing number of correct matches. Secondly, also inspired by CorresNeRF, we implemented what we call the “two-way” MV loss:
$\mathcal{L}_{{MV}}=\lambda_{MV}\textbf{w}_{\mathbf{p}_{i},\mathbf{p}_{j}}\frac{\sum_{n\in\{i,j\}}f_{huber}[\mathbf{x}({\mathbf{p}_{n}})-\mathbf{x}_{gt}({\mathbf{p}_{n})}]}{2}$
Compared to the Vanilla MV loss, the two-way MV loss additionally starts with known matches in pose $\mathbf{p}_{j}$. They are similarly projected to 3D, and back-projected onto the view at pose $\mathbf{p}_{i}$. This provides predicted matches $\mathbf{x}({\mathbf{p}_{i}})$, then compared with ground-truths $\mathbf{x}_{gt}({\mathbf{p}_{i}})$. Intuitively, the two-way MV loss can provide more rigid geometric supervisions, as in one iteration both directions are considered.
Implementation Details Again, to compare the DW and MV losses fairly, we simply replaced the DW loss with the MV loss in Algorithm 1, and kept other hyperparameters intact, except that the weight $\lambda_{MV}$ was set to 0.001. In each iteration, we sampled a pair of known views $\{\mathbf{p}_{i},\mathbf{p}_{j}\}$, and calculated the MV loss. We used a trained OmniGlue [OmniGlue] for predicting matches with good generalization. To better ensure matches quality, we filtered out correspondences with confidences $\textbf{w}_{\mathbf{p}_{i},\mathbf{p}_{j}}<0.5$. Figure 9 presents an example of matches from the “fern” scene. Even with confidence filtering, due to domain shifts between training and testing, incorrect matches still persist. This is a common drawback of correspondence methods [33].
As presented in Table 4, multi-view supervision degrades performance. While the method has consistently worked on slower-converging NeRF, we believe INGP’s aggressive optimization dynamics will amplify supervision of incorrect matches, leading to sub-optimal results than our frequency supervision, which does not require pre-trained weights as priors. We hope our results encourage fellow researchers to re-adapt and re-think correspondence methods for fast-converging representations like INGP and 3DGS [15].
11 Frequency supervision in novel poses
It can be advantageous to perform supervisions on novel poses [11]. In the end, the task is novel view synthesis. Therefore, we tried extending the DW loss to novel poses:
$\mathcal{L}_{DW}=\sum_{sb\in\{LL,LH,HL,HH\}}\lambda_{sb}||\mathbf{I}^{sb}(\mathbf{\hat{p}})-\tilde{\mathbf{I}}_{gt}^{sb}(\mathbf{\hat{p}})||_{2}^{2}$
where $\mathbf{\hat{p}}$ is an interpolated view from a sampled pair of known poses $\{\mathbf{p}_{0},\mathbf{p}_{1}\}$, similar to our implementation of DietNeRF. $\tilde{\mathbf{I}}_{gt}^{sb}(\mathbf{\hat{p}})$ is the pseudo ground-truth sub-band at novel view $\hat{\mathbf{p}}$. Inspired by GeCoNeRF [17] and SfMLearner [SfMLearner], we obtained $\tilde{\mathbf{I}}_{gt}^{sb}(\mathbf{\hat{p}})$ as follows:
$\tilde{\mathbf{I}}_{gt}^{sb}(\mathbf{\hat{p}})=\texttt{DWT}[\tilde{\mathbf{I}}_{gt}(\mathbf{\hat{p}})]=\texttt{DWT}\{\texttt{sampler}[\mathbf{I}_{gt}(\mathbf{p}),\mathbf{x}_{\mathbf{\hat{p}}\rightarrow\mathbf{p}}]\}$
where $\texttt{DWT}\{\cdot\}$ is the Discrete Wavelet Transform, operating on $\tilde{\mathbf{I}}_{gt}(\mathbf{\hat{p}})$. The $\texttt{sampler}[\cdot,\cdot]$ applies bilinear interpolation to retrieve the warped pixel values from the actual ground-truth view at known pose $\mathbf{p}$, $\mathbf{I}_{gt}(\mathbf{p})$, at the projected coordinates $\mathbf{x}_{\mathbf{\hat{p}}\rightarrow\mathbf{p}}$:
$\mathbf{x}_{\mathbf{\hat{p}}\rightarrow\mathbf{p}}=\mathbf{K}\mathbf{T}_{\mathbf{\hat{p}}\rightarrow\mathbf{p}}\mathbf{D}(\mathbf{\hat{p}})\mathbf{K}^{-1}\mathbf{x}_{\mathbf{\hat{p}}}$
where $\mathbf{K}$ is the known camera intrinsic matrix, $\mathbf{T}_{\mathbf{\hat{p}}\rightarrow\mathbf{p}}$ is the camera transformation matrix from the novel to known poses, $\mathbf{D}(\mathbf{\hat{p}})$ is the predicted depth map at the novel pose (naturally available as part of the rendering process) and $\mathbf{x}_{\mathbf{\hat{p}}}$ are pixel coordinates in the novel pose. Figure 10 shows a visual example of this warping process.
Figure 10: Differentiable image warping: (a) rendered view at novel pose $\mathbf{\hat{p}}$, b) its pseudo ground-truth view, which is warped from c), actual ground-truth view at known pose $\mathbf{p}$. The warping allows supervision in novel poses.
Implementation Details We again replace the Vanilla DW loss with the “Novel” DW loss in Algorithm 1, and kept all other hyperparameters intact. The warping process will introduce black regions due to out-of-bounds sampling, but we simply did not consider them in the loss calculation.
Table LABEL:tab:wavelet_abs_novel shows the results. The DW loss, operating on novel poses, still outperforms the baseline, but performance is worse than operating on known poses. We hypothesize this is because warping errors corrupt the frequency supervision, and because out-of-bounds points reduce information for supervision. Future works could explore integrating the semantic (which works well at novel poses) and frequency supervision (which does not). For example, if an object is what it is regardless of viewing direction, so are its low and middle frequencies (like edges). This is unlike high frequencies, which are much more vulnerable to changes in viewing directions. This could be an interesting direction of frequency supervision. In the meantime, this experiment shows that the DW loss does not require operating on novel poses, thereby maintaining its ease of implementation.
Table 7: Ablation on DW loss in novel poses
Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
CombiNeRF (baseline) 20.12 0.676 0.197
DWTNeRF (w/ known dw.) 20.38 0.677 0.193
DWTNeRF (w/ novel dw.) 20.22 0.669 0.194
Figure 11: Multi-level decompositions. The top row is for the ground-truth view, the bottom for the rendered view. From left to right: a) original image, b) 1-level decomposition and c) 2-level. Moving from b) to c), only the LL sub-band is decomposed further, while the remaining sub-bands are kept intact.
12 Frequency decomposition levels
Table 5 shows results for 1-level decomposition, where the DWT is applied once to provide the sub-bands. However, we can take the resulting LL sub-band and similarly decompose it into another set of 4 sub-bands:
$\mathbf{I}^{LL}_{2}=\mathbf{L}\mathbf{I}^{LL}\mathbf{L}^{T};\mathbf{I}^{LH}_{2}=\mathbf{H}\mathbf{I}^{LL}\mathbf{L}^{T}$
$\mathbf{I}^{HL}_{2}=\mathbf{L}\mathbf{I}^{LL}\mathbf{H}^{T};\mathbf{I}^{HH}_{2}=\mathbf{H}\mathbf{I}^{LL}\mathbf{H}^{T}$
where $\mathbf{I}^{(\cdot)}_{2}$ represent the 2-level decomposition of the sub-bands. $\mathbf{I}^{LL}=\mathbf{L}\mathbf{I}\mathbf{L}^{T}$ is the LL sub-band from 1-level decomposition. The remaining 1-level sub-bands are left intact. By further decomposing the LL sub-band, we can better enable the multi-resolution analysis characteristic of the DWT. Theoretically, this allows the representation of the scene at progressively coarser resolutions, enabling a hierarchical understanding of the scene’s structure. However, this is hard to implement. 1-level decomposed sub-bands are 4 times smaller than the original image. 2-level sub-bands are 8 times smaller, and so on. To ensure the sub-bands are still visually meaningful, and thus provide useful supervision, the original image has to be larger. As explained, this requires sampling a higher number of rays, making the method computationally expensive. For now, we conducted a simple experiment that used the 2-level decomposed sub-bands instead of 1-level in Algorithm 1. To keep the method computationally simple, we retained the patch-based sample batch size $|\mathcal{R}_{P}|$ used in the 1-level experiment. Figure 11 shows the decomposed scenes. 2-level decomposition loses finer details, but pushes the details to the forefront, so it’s worth investigating. Table LABEL:tab:wavelet_abs_decomp_lvl shows the results. While using 2-level still outperforms the baseline, the results are worse than 1-level. This confirms our intuition that the DW loss needs visually meaningful features. However, future works could introduce an efficient procedure to allow a multi-level, hierarchical DW loss that is more in line with multi-resolution analysis.
Table 8: Ablation on wavelet decomposition levels
Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
CombiNeRF (baseline) 20.12 0.676 0.197
DWTNeRF (+dw. / haar - 1-lvl) 20.25 0.677 0.193
DWTNeRF (+dw. / haar - 2-lvl) 20.21 0.677 0.194
13 Model methods

### v3

@@ 5.2 Qualitative Results
Method Venue PSNR $(\uparrow)$ SSIM $(\uparrow)$ LPIPS $(\downarrow)$
@@ 5.3 Quantitative Results
On LLFF. We compared DWTNeRF against recent few-shot works published at top venues. Regularization-based methods include InfoNeRF [16], FreeNeRF [39] and CombiNeRF [2]. Prior-based methods include DietNeRF [11], RegNeRF [23], DiffusioNeRF [37] and GeCoNeRF [17]. Model-based method includes mi-MLP [46]. Table 1 shows the quantitative results. Methods with the symbol $\dagger$ were reproduced on our machine, keeping all default settings. Unavailable metrics are denoted “-”. The best-performing metrics under each few-shot condition are bolded. In the 2-shot case, DWTNeRF outperforms CombiNeRF by 1.6% in PSNR, 0.84% in SSIM and 4.17% in LPIPS. The effects of our DWTNeRF are still strongly seen with 3-shot, but generally fade with higher views, suggesting its high compatibility with extreme few-shot conditions.
On NeRF-Synthetic. Table 2 shows the quantitative results. In the 4-shot case, DWTNeRF outperforms CombiNeRF by 0.5% in PSNR, 0.1% in SSIM and 0.9% in LPIPS.
@@ 5.4 Ablation Studies
On main modules. Those are the DW loss and the cross-banch interactions. Table 3 shows the effects of each module. The second row removes the frequency masking from CombiNeRF, given its incompatibility with INGP. This provides improvements in all metrics. The third row is our DWTNeRF, but with the DW loss only. Starting from the fourth row, we examine the effects of using multi-head attention in cross-branch interactions. The fourth row is DWTNeRF with only input-level attention. The fifth row only has output-level attention. The sixth row combines both levels. This produces stronger results over the previous two. The final row combines both DW loss and cross-branch interactions, which provides the strongest results.
@@
Table 4: Comparison against model-based methods, LLFF dataset [20]
@@
On model-based methods We implemented two model-based approaches on top of CombiNeRF: “Residual Connections” (Figure 6a) and “Element-wise Cross-branch Interactions” (Figure 6b) - both introduced by mi-MLP [46]. Table 4 shows the results. The second row introduces residuals. The third row is the same, but we added one MLP layer, hoping to better highlight the residuals’ effects. However, this degraded performances noticeably, highlighting the vulnerability of INGP to architectural changes. The fourth row shows mi-MLP’s element-wise cross-branch interactions. Only our version of interactions, based on attention, is competitive with the baseline, and exceeds in PSNR and LPIPS. For more implementation details, please refer to Section 9 of Appendix.
Figure 6: mi-MLP’s [46] model-based methods: “Residual Connections” (a) and “Element-wise Cross-branch Interactions” (b)
On the choice of wavelets. We compared the effects of different Daubechies wavelets [6] on the DW loss. We started with the Haar ($n=1$) and then higher-order wavelets, up to the db3 ($n=3$). Generally, higher-order wavelets have more supports, and spreads the signal energy more evenly across sub-bands. Section 8 of Appendix shows the exact wavelet coefficients. Figure 7 compares scene decomposition using the Haar versus db2 ($n=2$). Compared to the ground-truths, the rendered sub-bands are much noisier, highlighting the purpose of the DW loss. The db2 provides finer details. However, table 5 shows no significant differences between the wavelets. This confirms our intuition that low-frequency supervision is sufficient.
@@
Figure 7: Decomposition using the Haar (a & b) and db2 (c). We only show the LL (top) and LH (bottom) sub-bands. Column a) is the ground-truth $L$, much less noiser than rendered $L$ at column b). Column c) uses higher-order wavelet, and shows finer details.
@@ 6 Conclusion
Limitations & Future Works. Firstly, the DWT does not disentangle high-frequency components. For example, besides fine structural details, high frequencies can come from specularity and other light-dependent effects. Future works might explore scene decomposition at such further levels. Secondly, since the DW loss is used as an auxiliary loss alongside the photometric MSE loss, which compares pixel-wise differences regardless of frequencies, we believe some high-frequency overfitting still exists. Our DW loss emphasizes low frequencies, but not eliminate high frequencies in early training. Future works could tackle this problem, but interfering with the photometric loss can be tricky. We can also investigate the applicability of our methods to other representations, like SDF [3, 18] or 3DGS [15].
@@ 7 More Visualizations
We provide more qualitative results on the 3-shot LLFF and 4-shot NeRF-Synthetic benchmarks at Figures 10 and 9. There are 6 scenes in total. In each scene, we would like to direct attention towards the regions enclosed by the red box, which highlights the effects of cross-branch interactions (b $\rightarrow$ c) and DW loss (c $\rightarrow$ d). They provide finer details, preserve better structures and reduce random “hallucinated” points in 3D space.
@@ 8 Wavelet Coefficients
[…]
@@
[…]
@@
[…]
@@
Figure 8: Frequency-based scene decompositions with different wavelets. From top to bottom: the $L$, LH, HL and $H$ sub-bands. The first column (a) shows the sub-bands of ground-truth views using the Haar. The second column (b) is the same, but for rendered views, which are much noisier. The rest show sub-bands of rendered views using the db2 (c) and db3 (d). Higher-order wavelets provide more fine-grained details, but our experiments showed low-frequency supervision with coarse details is sufficient.
9 Model-based methods